Find dy/dx and d²y/dx²;


x=√t, y=2t+4 ; t=1

1
Expert's answer
2021-12-01T16:29:38-0500

"x = \sqrt t \Rightarrow \frac{{dx}}{{dt}} = \frac{1}{{2\sqrt t }}"

"y = 2t + 4 \Rightarrow \frac{{dy}}{{dt}} = 2"

Then

"\frac{{dy}}{{dx}} = \frac{{\frac{{dy}}{{dt}}}}{{\frac{{dx}}{{dt}}}} = \frac{2}{{\frac{1}{{2\sqrt t }}}} = 4\sqrt t \Rightarrow {\left. {\frac{{dy}}{{dx}}} \right|_{t = 1}} = 4"

"\frac{d}{{dt}}\left( {\frac{{dy}}{{dx}}} \right) = \frac{4}{{2\sqrt t }} = \frac{2}{{\sqrt t }}"

Then

"\frac{{{d^2}y}}{{d{x^2}}} = \frac{{\frac{d}{{dt}}\left( {\frac{{dy}}{{dx}}} \right)}}{{\frac{{dx}}{{dt}}}} = \frac{{\frac{2}{{\sqrt t }}}}{{\frac{1}{{2\sqrt t }}}} = \frac{{4\sqrt t }}{{\sqrt t }} = 4 \Rightarrow {\left. {\frac{{{d^2}y}}{{d{x^2}}}} \right|_{t = 1}} = 4"

Answer: "{\left. {\frac{{dy}}{{dx}}} \right|_{t = 1}} = 4" , "{\left. {\frac{{{d^2}y}}{{d{x^2}}}} \right|_{t = 1}} = 4"